n_TOF cooling station gas releases improvements

2011 n_TOF facility status
E. BERTHOUMIEUX ON BEHALF THE N_TOF
COLLABORATION
Outlines
 Proton status
 Neutron beam line status
 Status of the 2010 gas release problem
 Status of the moderator & cooling circuit
 Remaining RP issues
 Foreseen / possible facility upgrade



RP related
Access related
In the neutron beam line
 Web status
 DAQ update
Proton status
 Reminder: proton is read by a new transformer since
beginning of this year to solve the saturation problem
 For the 1st
time n_TOF
reached is
requested
intensity in
2011: 1.64 1019
 n_TOF Ph2
target received
more beam
than old one
(>3.4 1019 p)
Neutron beam line status
 Reminder: First collimator not (and can not be)
aligned.
 Small leakage at the filter station level (probably
with the C filter)
 Several general electricity cuts occurred this year
causing sometimes difficulties to restart the
pumping, the magnet or the DAQ
 Sweeping magnet switched off due to a “current
balance” alarm. No real causes found by the expert.
 EM & grounding problems in the WSTA.
 Fences have been replaced to avoid intrusion in the
WSTA
n_TOF inside ISR-8
n_TOF ventilation
~30 m2
n_TOF cooling
station ~70 m2
2010 gas released problem (EDMS 1147942)
 Maximum personnel exposure in 2010 due to the
n_TOF cooling station
External photon
dose rate
meas.
• Estimation of
relative
contribution of
the skin and
inhalation dose
• n_TOF operating
conditions (beam
and degassing)
• Occupation factor
Modifications on moderator circuit (reminder)
 Cationic to mixed bed (previously saturated in enriched
boron thanks to the bench test) resins.

No raise in conductivity seen since the filling of the circuit
 Moderator circuit was not designed to switch twice a year
from demineralized to borated water



New expansion vessel to fill and check the water level in the
moderator circuit
New valves and pipes to empty the circuit.
Filling/emptying only require now valves manipulations.
 Circuit was filled with enriched boric acid (not the same
as last year!!!) the 1st of march and boron (~1 kg) was
added in the “pot à bore” the 8th of march
Moderator station status
 Moderator circuit working at the saturation limit (filled
with water containing 41.8 g/l of boric acid)


Some difficulties to have a stable regime in the borated water
circulation at start-up.
Colder chilling water (7±2 ºC) and higher chilling water flow inside
the “pot à bore” to ensure that it remains the coolest point of the
moderator circuit
 Conductivity very stable up to the end of run (~57 µS)

Since 23 nov 2011 conductivity raised to 62 µS
 A water leakage (~1.5 l loss in 6 months) was found at the
TT9 sensor position.

Fixed but still under surveillance
 Very high tritium content (~17 MBq/l)

Borated water will be replaced during this winter shutdown
Cooling station status
 Shielding has been reinforced during 2010 shutdown
 Still not sufficient (see RP issues later)
 Leakage of the degassing circuit solved
 This circuit is now a radioactive source!
 Problem with the oxygen content measurement
 Value read by the station is wrong: shift of >0.04 mg/l
compared to the value directly read on the apparatus.
 Threshold have been modified to take care of this shift.
 Action to be done during next shutdown to solve the problem
RP issues (from J. Vollaire)
 Dose rate in the vicinity of the cooling station can be
still above the CERN limit for a non exposed area


Blocks added
(31/08/2011) to
overcome this issue
2 workplaces


Blocks added here
Center
Right
Detector at the center
Remaining RP issues (from J. Vollaire)
Before
raw
mSv/h
Before
norm. to
alarm mSv/h
After
raw
mSv/h
After norm.
to alarm
mSv/h
Reduction
PMI
3200
6000
2958
6000
FH40
(right)
0.34
0.63
0.27
0.54
15 %
FH40
(center)
0.29
0.54
0.25
0.51
6%
 Results are normalized to alarm A level on
PMITOF03 (= 6 mSv/h alarm B = 7 mSv/h)
 Effect seems visible but is not sufficient to declassify
the area with the current nominal intensity and
alarm levels
Summary of known problems in ISR8
Dose > 0.5 µSv/h
RP related possible facility upgrade (ISR8)
 Remaining dose at workplace area assigned to skyshine
effect.
 Dose in front of cooling station due to combination of
skyshine effect and degassing duct.
Foreseen/possible upgrades
 Permanent RP monitoring at workplace area (RP
action already started)
 Send degassing duct directly to primary area

The initial idea to just replace existing duct has been abandoned
 Roof (total or partial) over the station
Simple FLUKA model (J. Vollaire)
Shielded
Un-Shielded
Roof : 5 cm steel
Shielded case has
lower side walls : 2 x
80 cm
Conclusions
 Reduction of the
dose rate from a
factor 10-20 in the
TE workshop area
 Reduction of factor
~2 in front (second
wall not in the
simulation) due to
opening
Access system in primary area will be revised
(EDMS 1145465, 1145476 & 990214)
 PS complex access system will be completely revised
during the long shutdown
 GS/ASE
mandate will be
focused on
access only


COOLING
TARGET
Ventilation of
primary area will
be part of the
access system
Cooling station
will no longer be
interlocked with
beam!!!
PORTE FIN DE ZONE GRAND
LARGEUR (2600 mm) POUR
PASSAGE ENGINS
PASSAGE
18 avril 2011
upgrades in the neutron beam line
 Filter station
 Solve the leakage problem
 Add/replace some filters?
 1st collimator alignment?
 Magnet control (~30 years old, lack of spares pieces)
 Grounding improvement in WSTA
 EM decoupling of the beam pipe nearby the WSTA?
 Improve alignment in WSTA?
n_TOF web status
 Old web site: pceet075.cern.ch
 No longer maintained: migration needed!
 TWiki website: cern.ch/ntof (default website)
 Need to register in TWiki group NtofGroup to access and modify some parts of
the site
 DAQ website: ntofdaq.cern.ch
 Need CERN credentials + ntof-runs e-group membership
 Pictures gallery: ntofweb.cern.ch/gallery2
 No longer maintained: moved to DAQ website
 Incoming public website: cern.ch/ntof-exp
 Need volonteers to build/transfer infos from old website
 n_TOF on indico: indico.cern.ch
 To be used for any n_TOF related meeting
 n_TOF section on CERN doc. server: cdsweb.cern.ch
 To publish any note/paper related to n_TOF
A few word about DAQ
Hardware
 Got 4 additional channels (DC282 module) from
Zagreb

Need to add memory inside
 Bought 4 2nd hand DC282 modules (16 channels)
 Need to add another stream in the DAQ to be able to use it